# An Optimization Framework to Enforce Multi-View Consistency for Texturing 3D Meshes Using Pre-Trained Text-to-Image Models

### v1

@@ 4.4 Stage IV: Texture Stitching
The final stage of our approach takes the aligned RGB-D images $I_{i}^{\star}=\mathcal{D}_{i}^{\star}(I_{i}),1\leq i\leq k$ as input and outputs a complete textured 3D model. This is formulated as decomposing the mesh faces $\mathcal{F}$ of the input mesh $\mathcal{M}$ into sets $k$: $\mathcal{F}=\mathcal{F}_{1}\cup\cdots\mathcal{F}_{k}$ where $\mathcal{F}_{i}$ is associated with $I_{i}^{\star}$. We formulate this procedure as a joint labeling problem. Specifically, we associate each face $f\in\mathcal{F}$ with a random variable $z_{f}\in\{1,\cdots,k\}$ that specifies the set of faces $\mathcal{F}_{z_{f}}$ to which it belongs. Let =$\mathcal{E}=\{(f,f^{\prime})\}$ collect all pairs of adjacent faces. We optimize $z_{f},f\in\mathcal{F}$ jointly by minimizing the following objective function:
@@
where $\phi_{f}(z_{f})$ is the cost of label $z_{f}$, and $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})$ is the consistency between $z_{f}$ and $z_{f^{\prime}}$. In this paper, we define $\phi_{f}(z_{f})$ using a variant of (5), that is, a label has a small cost if it is more consistent with other images. Furthermore, $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})=0$ if $z_{f}=z_{f^{\prime}}$. Otherwise, its value depends on the color difference between $z_{f}$ and $z_{f^{\prime}}$. Details are deferred to the supp. material.
We solve (9) using the OPENGM2 [16] implementation of TRWS [19]. Figure 7 (a) shows an example of optimized cuts between aligned images.
After obtaining $\mathcal{F}_{i},1\leq i\leq k$, we use the corresponding cut between each pair of overlapping images $I_{i}$ and $I_{j}$ to recompute $O_{ij}$ and $O_{ji}$. This is achieved by dilating the cut in each $I_{i}$ by $l$ pixels ($l=20$) and taking the union of the dilated regions to define $O_{ij}$ and $O_{ji}$. After that, we rerun the view alignment to generate the inputs for texture stitching. This alternating procedure is iterated until the final output becomes steady (typically after 2-3 iterations). Figure 7 (b) shows that this alternate procedure can enhance visual consistency along the cuts.
@@ 5 Experimental Results
We begin with the experimental setup in Section 5.1. We then present the experimental results in Section 5.2. Finally, we describe an ablation study in Section 5.3.
@@ 5.1 Experimental Setup
Dataset. We collect 20 3D models from existing text-to-textured-mesh methods and 300 more models from different categories on the GObjaverse dataset [32].
@@ 5.2 Analysis of Results
FID(↓) 67.85 57.29 51.34 48.26
As shown in Table 1, the quantitative results are consistent with the relative improvement in visual performance.
This result quantitatively justifies our design that combines view generation, selection, alignment, and view stitching.
@@ 5.3 Ablation Study
This section describes an ablation study of our approach. Qualitative and quantitative results are given in Figure 11 and Table 1 respectively.
Figure 11: Visualization of Ablation Study
No view selection. View alignment is a critical component of our approach. Without view selection, the FID score increases by 40
No view alignment. View alignment is another critical component of our approach. Without view alignment, the FID score increases by 18
No alternating optimization. Without alternating optimization, the performance of our approach drops both qualitatively and quantitatively. The FID score increases by 6
@@ 7 Conclusions and Future Work
Supplementary Material
8 Multi-View Consistent Initialization
We utilize the multi-view depth-conditioned albedo diffusion model from RichDreamer  [32] for the initialization. The rendered depth maps are concatenated with the latent features to serve as input for the UNet denoiser [33]. We only sample 8 images around the underlying 3D model instead of an overcomplete set, and we unproject all images to get an averaged RGB UV map as initialization. The invisible regions on the UV map are extracted by pixel detection and inpainted [46] by the region neighborhood. The inpainted UV map is then decoded to latent space and noised to the specified diffusion step. Although the sampled images have fewer details and are averaged in UV space, this is sufficient for initialization [27].
9 Texture Stitching details
In this section, we provide details on the formulations of $\phi_{f}(z_{f})$ and $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})$.
Formulation of $\phi_{f}(z_{f})$. With $\mathcal{C}_{f,i}=\{\boldsymbol{c}\}$ we denote all colors of pixels of image $I_{i}$ associated with face $f$. Let $\mathcal{C}_{f}=\cup_{1\leq i\leq k}$ collect all the colors of the pixels associated with the face $f$. We perform mean sift clustering among $\mathcal{C}_{f}$. With $\boldsymbol{c}_{f}$ and $\sigma_{f}$ we denote the center of the cluster and the associated variance, respectively. We then define
$\phi_{f}(z_{f})=\left\{\begin{array}[]{cc}+\infty&\mathcal{C}_{f,z_{f}}=\emptyset\\ \frac{1}{\mathcal{C}_{f,z_{f}}}\sum\limits_{\boldsymbol{c}\in\mathcal{C}_{f,z_{f}}}\frac{\|\boldsymbol{c}-\boldsymbol{c}_{f}\|^{2}}{2\sigma_{f}^{2}}&\textup{otherwise}\end{array}\right.\ .$ (10)
Formulation of $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})$. Denote $\mathcal{P}_{f,i}$ as the set of pixels in $I_{i}$ that belong to face $f$. If $\mathcal{P}_{f,i}\neq\emptyset$, we compute $\boldsymbol{f}_{f,i}=(\overline{\boldsymbol{c}}_{f,i},\mu\overline{\boldsymbol{d}}^{\textup{SIFT}}_{f,i}$, where $\overline{\boldsymbol{c}}_{f,i}$ is the average pixel color among $\mathcal{P}_{f,i}$ and $\overline{\boldsymbol{d}}^{\textup{SIFT}}_{f,i})$ is the average SIFT pixel descriptor among $\mathcal{P}_{f,i}$. We set $\mu=1$ in this paper. Note that instead of merely using color differences, we observe that incorporating SIFT helps place cuts among textureless regions.
Our definition of $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})$ differs from that of  [48] in the sense that we take advantage of the consistency of the images, in contrast to simply using color differences. Specifically,
  • When $\mathcal{P}_{f,z_{f}}=\emptyset$ or $\mathcal{P}_{f^{\prime},z_{f^{\prime}}}=\emptyset$,
    $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})=+\infty.$
  • When $\mathcal{P}_{f,z_{f^{\prime}}}=\emptyset$ and $\mathcal{P}_{f^{\prime},z_{f}}=\emptyset$,
    $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})=\|\boldsymbol{f}_{f,z_{f}}-\boldsymbol{f}_{f^{\prime},z_{f^{\prime}}}\|^{2}.$ (11)
  • When $\mathcal{P}_{f,z_{f^{\prime}}}\neq\emptyset$ and $\mathcal{P}_{f^{\prime},z_{f}}=\emptyset$,
    $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})=\|\boldsymbol{f}_{f,z_{f}}-\boldsymbol{f}_{f,z_{f^{\prime}}}\|^{2}.$ (12)
  • When $\mathcal{P}_{f,z_{f^{\prime}}}=\emptyset$ and $\mathcal{P}_{f^{\prime},z_{f}}\neq\emptyset$,
    $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})=\|\boldsymbol{f}_{f,z_{f}}-\boldsymbol{f}_{f^{\prime},z_{f}}\|^{2}.$ (13)
  • When $\mathcal{P}_{f,z_{f^{\prime}}}\neq\emptyset$ and $\mathcal{P}_{f^{\prime},z_{f}}\neq\emptyset$,
    $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})=\frac{1}{2}\Big{(}\|\boldsymbol{f}_{f,z_{f}}-\boldsymbol{f}_{f^{\prime},z_{f}}\|^{2}+\|\boldsymbol{f}_{f,z_{f}}-\boldsymbol{f}_{f,z_{f^{\prime}}}\|^{2}\Big{)}.$ (14)
Note that we only compute the color difference in (11) when $z_{f}$ and $z_{f^{\prime}}$ are not available in either $f$ or $f^{\prime}$. Otherwise, we employ the color consistency in (12), (13), and (14) which better reveals the appearance continuity after stitching.
10 Additional Results
10.1 Additional Baselines
The latest non-open source methods TexFusion [1] and Decorate3D [11] are compared with our method in Figure 12 and Figure 13, from where we see that our results are more natural in color and have clearer details.
10.2 More Visual Results
Figure 14 shows the results of different prompts on the same mesh. We present more visual results on objects from different categories in GObjaverse dataset [32] in Figures 15 to 17. Each textured mesh is rendered to 8 views. A full list of object indexes and prompts can be found in Table 2.
Figure 12: Qualitative comparison with TexFusion: clearer details.
Figure 13: Qualitative comparison with Decorate3D: more natural in color.
Figure 14: Different prompts for the same mesh.
[Uncaptioned image][Uncaptioned image]
Figure 15: Gallery of textured meshes.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 16: Gallery of textured meshes.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 17: Gallery of textured meshes.
Object Index Prompt
0002c6eafa154e8bb08ebafb715a8d46 Birthday cake
000f88bb21164319ae797d315be6bc0e Watch Model With Stand
0023717f4f564cc99f4ded70db04f590 Rusty kerosene lamp
0023b3edbc114be188ca9d8f729dfaaf a vintage car
0025c5e2333949feb1db259d4ff08dbe A wooden bird house
00286954e2d54db8bc7832cc8682b6ff Furniture Bed
002e02c30121465c8a01bcb83b584ea5 A pair of blue jeans
003199cc6ff2410cb2d8e6f8a9cbb163 a Portugal Fire Hydrant
0032696f5871429fbd0549d9628f812c A Christmas present
0033322379a24798a6875a5cb2de54f5 A raspberry
0046f208ef8d4988ba7bb9d297f29ec7 An old Dutch windmill building
004fb4d72f6c4e55a15b9025a868d1a3 a roast chicken
0056880681c044cb9fe815a9eed0425d A rhino
00602ef508784e5384665aacaaf1f3a0 A handle saw
0064add4992b426cb2f862e5875ebf6d A pink donut
0083fa5f10a442408e0f3f88df19c8ad An ancient shield
0087dc01648d4cc792a7d1e49848b825 Stone horse head sculpture
00978a128283411582590096643ec101 A pale yellow seahorse
00b267b43669422cbb4ec3a4e9b1c16e A pair of sunglasses
00d56831f9bc49f9a668f418c1af7558 A dancing brown dog in a white shirt, wearing dark sunglasses
010b9ece8a3a49e3b73be0b3cd02c720 A brown vase with red roses
013c3a1d945a4336a87f889c3d4c25b1 Precision Sniper Rifle, CSGO AWP DRAGON LORE
015777939fc3429ba4b5343be9d51ffa A kid’s bike
017fe235577b4083ab32c2b7949ba022 A platinum ring studded with precious stones
Table 2: The object indexes and the corresponding text prompts.

### v2

@@ 4.4 Stage IV: Texture Stitching
The final stage of our approach takes the aligned RGB-D images $I_{i}^{\star}=\mathcal{D}_{i}^{\star}(I_{i}),1\leq i\leq k$ as input and outputs a complete textured 3D model. This is formulated as decomposing the mesh faces $\mathcal{F}$ of the input mesh $\mathcal{M}$ into sets $k$: $\mathcal{F}=\mathcal{F}_{1}\cup\cdots\mathcal{F}_{k}$ where $\mathcal{F}_{i}$ is associated with $I_{i}^{\star}$. We formulate this procedure as a joint labeling problem. Specifically, we associate each face $f\in\mathcal{F}$ with a random variable $z_{f}\in\{1,\cdots,k\}$ that specifies the set of faces $\mathcal{F}_{z_{f}}$ to which it belongs. Let =$\mathcal{E}=\{(f,f^{\prime})\}$ collect all pairs of adjacent faces. We optimize $z_{f}$ by minimizing the following objective function:
@@
where $\phi_{f}(z_{f})$ is the cost of label $z_{f}$, and $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})$ is the consistency between $z_{f}$ and $z_{f^{\prime}}$. In this paper, we define $\phi_{f}(z_{f})$ using a variant of Eq. 5, that is, a label has a small cost if it is more consistent with other images. Furthermore, $\phi_{ff^{\prime}}(z_{f},z_{f^{\prime}})=0$ if $z_{f}=z_{f^{\prime}}$. Otherwise, its value depends on the color difference between $z_{f}$ and $z_{f^{\prime}}$. Details are deferred to the supp. material.
We solve Eq. 9 using the OPENGM2 [17] implementation of TRWS [20]. Figure 6 (Left) shows an example of optimized cuts between aligned images.
After obtaining $\mathcal{F}_{i},1\leq i\leq k$, we use the corresponding cut between each pair of overlapping images $I_{i}$ and $I_{j}$ to recompute $O_{ij}$ and $O_{ji}$. This is achieved by dilating the cut in each $I_{i}$ by $l$ pixels ($l=20$) and taking the union of the dilated regions to define $O_{ij}$ and $O_{ji}$. After that, we rerun the view alignment to generate the inputs for texture stitching. This alternating procedure is iterated until the final output becomes steady (typically 2-3 iterations). Figure 6 (Right) shows that this alternate procedure can enhance visual consistency along the cuts.
@@ 5 Experimental Results
We begin with the experimental setup in Sec. 5.1. We then present the experimental results in Sec. 5.2. Finally, we describe an ablation study in Sec. 5.3.
@@ 5.1 Experimental Setup
Dataset. We collect 300 models from different categories in Objaverse [7]. We also collect 20 models used in other text-to-texture-mesh methods and models from generation approaches [32, 9] for qualitative comparison.
@@ 5.2 Analysis of Results
FID(↓) 41.59 35.24 31.55 29.26
As shown in Tab. 1, the quantitative results are consistent with the relative improvement in visual performance. The blue numbers mean the second best performance. This result quantitatively justifies our design that combines view generation, selection, alignment, and view stitching.
@@ 5.3 Ablation Study
This section describes an ablation study of our approach. Quantitative results are given in and Tab. 1 and a qualitative result is available in supp. material.
No view selection. View alignment is a critical component of our approach. Without view selection, the FID score increases by 42%. Our observation is that without view selection, the textured mesh has appearance artifacts from erroneously selected images. There are also regions where the texture appearance is discontinuous. The latter issue is caused by inconsistent image alignment, where input affects output, despite the fact that robust alignment is used.
No view alignment. View alignment is another critical component of our approach. Without view alignment, the FID score increases by 20%. In this case, there are discontinuities across the textured mesh. These effects are caused by local multiview inconsistencies among the selected views. This issue cannot be easily addressed by solving an MRF problem to identify the cuts across the most consistent regions.
No alternating optimization. Without alternating optimization, the performance of our approach drops both qualitatively and quantitatively. The FID score increases by 8%. As shown in Fig. 6(a), we see some discontinuities across the cut. These are caused by the view alignment, in which pairwise alignments take into account all overlapping regions, even though some regions disagree. With alternating optimization, view alignments tend to be more consistent along the cuts, leading to improved results after cutting.